OpenStudy (anonymous):

solve by the roots of polynomial equation 5x^2-30x+45=0

OpenStudy (bossvideogamer21):

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 30*x-45-(5*x^2)=0 Step by step solution : Step 1 : Raise x to the 2nd power Exponentiation : Equation at the end of step 1 : (30x - 45) - (5 • x2) = 0 Step 2 : Simplify 30x-45 - 5x2 Step 3 : Pulling out like terms : 2.1 Pull out like factors : -5x2 + 30x - 45 = -5 • (x2 - 6x + 9)